The Bush administration has two things on their minds--obtaining funding for the border wall, and reaching a trade agreement with China. The probability of congress approving the proposed border budget is 0.73, while acquiring a memorandum of understanding with China has probability 0.98. Assume that these two events are independent.

a). What is the probability that the administration achieves both goals, given that at least one is achieved?

b). Conditional on a successful trade agreement, what is the probability that both goals are achieved?
Answer #1

P(A) = 0.73
P(B) = 0.98

a)

P(A or B) = P(A) + P(B) - P(A)*P(B) = 0.73 + 0.98 - 0.73*0.98 = 0.9946

P(A and B) = 0.73*0.98 = 0.7154

Required probability = 0.7154/0.9946 = 0.7193

b)
required probability = 0.7154/0.98 = 0.73